How many liters of pure baking soda must be added to 200 liters of a 30% baking soda solution to get a 60% baking soda solution?
Answer by ewatrrr(24785)


Hi,
How many liters of pure baking soda must be added to get a 60% baking soda solution
Let x be the number of liters of pure baking soda to be added
.30*100L + 1.00*x = .60(100L + x)
.30*100L + 1.00*x = .60*100L + .60x
simplify and solve
x - .60x = .60*100L - .30*100L
.4x = .30*100L
x = 75L
checking our answer
30L + 75L = .60*175L = 105L